# EEG classification for visual brain decoding with spatio-temporal and transformer based paradigms

### v1

@@ Abstract
In this work, we delve into the EEG classification task in the domain of visual brain decoding via two frameworks, involving two different learning paradigms. Considering the spatio-temporal nature of EEG data, one of our frameworks is based on a CNN-BiLSTM model. The other involves a CNN-Transformer architecture which inherently involves the more versatile attention based learning paradigm. In both cases, a special 1D-CNN feature extraction module is used to generate the initial embeddings with 1D convolutions in the time and the EEG channel domains. Considering the EEG signals are noisy, non stationary and the discriminative features are even less clear (than in semantically structured data such as text or image), we also follow a window-based classification followed by majority voting during inference, to yield labels at a signal level. To illustrate how brain patterns correlate with different image classes, we visualize t-SNE plots of the BiLSTM embeddings alongside brain activation maps for the top 10 classes. These visualizations provide insightful revelations into the distinct neural signatures associated with each visual category, showcasing the BiLSTM’s capability to capture and represent the discriminative brain activity linked to visual stimuli. We demonstrate the performance of our approach on the updated EEG-Imagenet dataset with positive comparisons with state-of-the-art methods.
@@ 1 Introduction
Hybrid models combine features of CNNs, RNNs/LSTMs, and sometimes GCNs to leverage the strengths of each in processing EEG signals [11] and can capture both spatial and temporal features efficiently.These models are versatile and can be tailored for a wide range of EEG-based visual classification tasks, from basic stimulus categorization to more complex applications like emotion recognition or neurofeedback.
@@
2 Related Work
In our examination of prior research, we scrutinized various methodologies for classifying Electroencephalography (EEG) data, particularly in the context of visual tasks. Li et al. [2] developed SyncNet, a model that employs structured 1D convolution layers to adeptly extract information from both the time and frequency domains, subsequently classifying the data through the integrated use of 1D convolutional neural networks (CNNs). On another front, Lawhern et al. [1] introduced EEGNet, which leverages 2D CNNs across different dimensions of EEG data.
Spampinato et al. [8] introduced a pioneering method for automated visual classification, utilizing human brain signals. They employed a CNN-based regression model to navigate the complex landscape of the EEG manifold. This approach involved the use of Recurrent Neural Networks (RNNs) to process EEG data elicited by visual image stimuli, which were then classified into a representation learned from the EEG signals.
Li et al. [2] critiqued the methodology employed by Spampinato et al. [8], pointing out that their reported results were contingent on a block design approach that may not hold up under a rapid-event design process. Specifically, they observed that the division of training and testing sets in Spampinato et al.’s study ensured that every trial in the test sets was derived from a block that was well-represented in the corresponding training set. Li et al. further argued that this methodology inadvertently boosted classification accuracy by capturing the long-term brain activity linked to the blocks of trials, rather than accurately reflecting the brain’s response to individual class stimuli.
Palazzo et al. [6] revisited their earlier research to address the criticisms raised by Li et al. [2]. This introspection revealed that the original classification performance, which they had reported with an average accuracy of around 83%, was indeed inflated. By applying correct filters, particularly targeting the high-frequency gamma-band and more realistic settings, the accuracy of classification showed a marked difference: their method achieved nearly 20% accuracy, EEGNet [1] reported about 30%, and EEG-Channel Net [6] saw a jump to approximately 50% accuracy. This adjustment not only highlighted the significance of proper data filtering but also vindicated their block design approach for classification studies.
This discovery implies that the models developed in previous studies cannot be directly compared due to their reliance on unfiltered EEG data, while the filtered dataset was only made available in 2020. In this study, our analysis and comparative evaluations are exclusively focused on methodologies that have utilized the filtered version of the EEG-ImageNet dataset. This decision stems from our aim to ensure a consistent and fair comparison across different approaches using the same data and settings.
In light of this, Tao et al. [9] conducted a comprehensive comparison of all leading models using the newly filtered data, exploring different frequency ranges: 55-95 Hz (high gamma band), 14-70 Hz, and 5-95 Hz (covering nearly the full spectrum). Their innovative model, based on the GRUGate Transformer, demonstrated a notable 61% accuracy when tested with data filtered for the high gamma band but showed a reduced accuracy of 49% when using data across the full frequency range.
In the realm of existing models, Viral Parekh et. al [7] developed an EEG-based image annotation system utilizing Convolutional Neural Networks (CNNs). This system leverages CNNs to extract static, local information from EEG data at specific timestamps. In contrast, the Perceive Laboratory [8] explored the use of Recurrent Neural Networks (RNNs) for creating models capable of visual discrimination. RNNs excel at maintaining and deriving features from long-term sequential data, capturing patterns from past inputs. However, one limitation of RNNs is their tendency to overlook the dynamic relationships between past and future electrical activations in relation to the present moment.
Considering the above approaches, and in order to make further advances in this domain, the contributions of this work involve:
@@ 2 Related Work
  • Positive comparisons with various approaches, and visualization involving t-SNE and brain mappings to analyze the discrimination of learned embeddings.
@@ 4 Methodology
Figure 2: This is the architecture of our model consisting feature block, sequential block and classifier block. Feature extractor block contains CNN layers to find the spatial information across both channel and time stamp while sequential block learns sequential relations using either Bi-LSTM or Transformer. The classifier containing dense layer classifies the embedding provided by sequential block.
@@ 4.2 Feature Extraction Block
This block contains CNN to separately extract spatial and temporal features across EEG channels and time axis. It performs 1-D convolution for three layers. 1-D convolution not only contains less parameters as compared to 2-D convolution but also helps in extracting temporal features from each channel and spatial features across channels, explicitly. This provides more control over generating embeddings in temporal and channel information without intermixing them. This is well explained in [5] and [4]. The first two layers perform convolution along time (activation function = Relu, kernel size = (1,35) and stride = (1,2)) while last layer do convolution over channel (activation function = sigmoid, kernel size = (128,1) and stride = (1,1)). The convolution layers over time find features across different time stamps while convolution layer over channel finds features across all channels for each extracted feature of time stamp. Each convolution layer has 25 filters. The final output of feature extractor is reshaped to (30,25) and fed to sequential block.
@@ 4.3 Bi-LSTM and Transformer Blocks
EEG signal can be fed to LSTM in two ways: either channel[8] wise or all EEG channel vector per time step. In common LSTM architecture whole EEG signal per time step is fed to LSTM. In Stacked LSTM, multiple layer of LSTM are used. Output of first layer is given to next layer. The final output of last layer at the last time step is the encoding of EEG signal[3], which can further be used for classification. In the Bi-LSTM block, we have stacked Bi-LSTM to capture information in both forward and backward direction.
Figure 3: Architecture of stacked Bi-LSTM in sequential block which is fed with extracted features in both forward and reverse manner in order to learn sequence in both forward and backward direction. Present feature is represented as x(t) while previous feature is represented as x(t-1). The forward and backward output of last layer of Bi-LSTM is combined together to fed it to classifier block.
@@
The Transformer model, consisting of an encoder and decoder architecture for sequence to sequence mapping. The encoder processes the input data utilizing self-attention mechanisms to weigh the non-local importance among different parts of data. This processed information is then passed to the decoder, which also employs self-attention and additionally uses what is known as cross-attention to focus on relevant parts of the input sequence when generating new output sequence. This allows Transformers to efficiently handle sequences of data, capturing more complex relationships within data compared to recurrent layers.
The Architecture mentioned in [10] consist of encoder and decoder for sequence to sequence mapping. As our take only involves EEG sequence classification, we therefore use only the encoder part of the transformer. The encoder of the transformer for our approach finds non-local relations within the sequential feature embeddings provided by the CNN using self-attention. The proposed architecture as shown in Figure 4 has 1 encoder layer with 8 heads for multi-head attention and 2 dense layers in feed forward network.
@@ 6.1 Experimental Setup
For our experiment, the split of 80% training, 10% validation and 10% test data is used as done in [9]. Each split contains data of each class for each subject. EEG data of all subjects for single image of a particular class is also taken into single split (either training or validation or testing) for uniformity. While splitting data it was observed that 33rd class has irregular data among different subjects. To maintain regularity among subject and class data, the 33rd class is removed, this resulting in a total of 39 classes.
@@ 6.2 Results
The result of the architectures and comparison with other existing work is shown in Table 1. These are either from their respective papers or from [3], which reports a comparison of various approaches.
@@
Approach Accuracy
Stacked Bi-LSTM [6] [8] 0.22
[8] 0.26
SyncNet [2] 0.27
[6] 0.31
EEGNet [1] 0.32
EEG-ChannelNet [6] 0.36
GRU Gated Transformer [9] 0.46
SigTanh Gated Transformer [9] 0.49
EfficientNet + SVM [3] 0.64
Our Bi-LSTM approach 0.71
Our Transformer approach 0.59
We note that both our CNN-Bi-LSTM based method outperform all existing approaches and the CNN-Transformer based method also outperform the most. The Bi-LSTM approach even performs better over approach in [11] which involves converting EEG signals to gray-scale image.
Table 2: Comparision of our Transformer approach with other existing Transformer approaches
Approach Accuracy
GRU Gated Transformer 0.46
SigTanh Gated Transformer 0.49
Our Transformer approach 0.59
We note that, presently the CNN-Transformer method does not perform as well as the CNN-Bi-LSTM approach. We believe that a reason for the relatively low performance of Transformer based design could be due to smaller window length, since Transformers usually expects long duration data better learn non-local dependencies. Moreover, the architecture and training dynamics of Transformers might necessitate adjustments for optimal performance on EEG datasets. Factors such as the choice of hyper parameters, the design of the attention mechanism, and the amount of training data available could significantly impact the effectiveness of Transformers as this is an unconventional application domain. Hence, the unique challenges posed by EEG data require bespoke modifications to the standard Transformer model to fully leverage its capabilities for temporal feature extraction.
Having said that, from Table 2, we also note that the Transformer based framework in this work outperforms the other existing Transformer based approaches which uses a gated network (using either GRU Gate or SigTanh Gate) in place of the residual network of the encoder of Transformer. Thus, clearly this work highlights the potential using Transformer based architectures in this domain.
@@ 7.0.1 Acknowledgements
The work has been partly supported by Science and Engineering Research Board, Government of India.

### v2

# refrences.bib

@@ Abstract
In this work, we delve into the EEG classification task in the domain of visual brain decoding via two frameworks, involving two different learning paradigms. Considering the spatio-temporal nature of EEG data, one of our frameworks is based on a CNN-BiLSTM model. The other involves a CNN-Transformer architecture which inherently involves the more versatile attention based learning paradigm. In both cases, a special 1D-CNN feature extraction module is used to generate the initial embeddings with 1D convolutions in the time and the EEG channel domains. Considering the EEG signals are noisy, non stationary and the discriminative features are even less clear (than in semantically structured data such as text or image), we also follow a window-based classification followed by majority voting during inference, to yield labels at a signal level. To illustrate how brain patterns correlate with different image classes, we visualize t-SNE plots of the BiLSTM embeddings alongside brain activation maps for the top 10 classes. These visualizations provide insightful revelations into the distinct neural signatures associated with each visual category, showcasing the BiLSTM’s capability to capture and represent the discriminative brain activity linked to visual stimuli. We demonstrate the performance of our approach on the updated EEG-Imagenet dataset with positive comparisons with state-of-the-art methods. Our best performing model yields an accuracy of 71%, a significant improvement over various existing methods.
@@ 1 Introduction
Hybrid models combine features of CNNs, RNNs/LSTMs, and sometimes GCNs to leverage the strengths of each in processing EEG signals [zhang2020bi] and can capture both spatial and temporal features efficiently.These models are versatile and can be tailored for a wide range of EEG-based visual classification tasks, from basic stimulus categorization to more complex applications like emotion recognition or neurofeedback.
@@
The domain of visual brain decoding is relatively recent, the existing work in this area primarily employs either CNN architectures or sequential architectures such as Bi-LSTM and transformers. Our method, however, integrates both CNN and sequential architectures (Bi-LSTM and transformer), thereby utilizing both local and temporal features for classification. Our work demonstrates that such frameworks, while common to other areas, can benefit EEG brain decoding, considering that CNN filter embeddings help in learning features from noisy EEG data, followed by sequential and attention based architectures, which exploit natural temporal dependencies in EEG signals.
Thus, the contributions of this work involve:
@@
  • Positive comparisons with various approaches, and analysis involving t-SNE and brain mappings to analyze the discrimination of learned embeddings.
2 Related Work
Our classification task includes EEG signals which are sequence of electrical signals captured over time through electrodes. Traditional methods done on EEG such as spectrogram, continous wavelet transform are used to extract features from raw EEG [thodoroff2016learning], [tsinalis2016automatic]. Different classifiers such as k-nearest neighbor, multi layer perceptron and Support Vector Machine (SVM) are also used for classification [li2020perils]. After the introduction of neural networks, Convolutional Neural Network (CNN) [li2020perils], [craik2019deep] and Recurrent Convolutional Neural Network (RCNN) [bashivan2015learning] are also used for classification. People have also used Long Short Term Memory (LSTM) [li2020perils], [bozal2017personalized] and Recurrent Neural Network (RNN) [craik2019deep] for classification based on EEG.
For our task to classify EEG signals to one of 40 different classes of ImageNet, people have opted various methods to approach the problem. SyncNet [li2017targeting] and EEGNet [lawhern2018eegnet] are two popular models based on CNN architecture. Li et al. [li2017targeting] developed SyncNet, a model that employs structured 1D convolution layers to adeptly extract information from both the time and frequency domains, subsequently classifying the data through the integrated use of 1D convolutional neural networks (CNNs). On another front, Lawhern et al. [lawhern2018eegnet] introduced EEGNet, which leverages 2D CNNs across different dimensions of EEG data. [zhang2020bi] has converted EEG signals to grayscale images then performed the classification task by extracting features from it. Many papers have used Siamese network [palazzo2020decoding]and similar approach [spampinato2017deep], [shimizu2022improving] to first bring the feature space of EEG and its corresponding image to same latent space using CNN or RNN and then use classifier for classification of EEG signals.
Since EEG is a sequence, an architecture rich in storing sequential data should be considered. RNNs are prone to vanish gradient for large sequences. Therefore, we have opted transformer and Bidirectional Long Short Term Memory (Bi-LSTM).
There are different existing methods using this approach. [tao2021gated] has changed the residual connection in transformer after attention and feed forward network with GRU or SigTanh calling it as Gated Transformer. [xie2022transformer] has used transformer with CNN in different combination to capture different information such as channel, temporal and fusion of both with different types of positional embeddings. [fares2019eeg] used Bi-LSTM to learn features and then select them using independent component analysis (ICA) for classifying using SVM.
Li et al. [li2017targeting] also critiqued the methodology employed by Spampinato et al. [spampinato2017deep], pointing out that their reported results were contingent on a block design approach that may not hold up under a rapid-event design process. Specifically, they observed that the division of training and testing sets in Spampinato et al.’s study ensured that every trial in the test sets was derived from a block that was well-represented in the corresponding training set. Li et al. further argued that this methodology inadvertently boosted classification accuracy by capturing the long-term brain activity linked to the blocks of trials, rather than accurately reflecting the brain’s response to individual class stimuli.
Palazzo et al. [palazzo2020decoding] revisited their earlier research to address the criticisms raised by Li et al. [li2017targeting]. This introspection revealed that the original classification performance, which they had reported with an average accuracy of around 83%, was indeed inflated.
By applying correct filters, particularly targeting the high-frequency gamma-band and more realistic settings, the accuracy of classification showed a marked difference: their method achieved nearly 20% accuracy. This adjustment not only highlighted the significance of proper data filtering but also vindicated their block design approach for classification studies.
This discovery implies that the models developed in previous studies cannot be directly compared due to their reliance on unfiltered EEG data, while the filtered dataset was only made available in 2020. In this study, our analysis and comparative evaluations are exclusively focused on methodologies that have utilized the filtered version of the EEG-ImageNet dataset. [mishra2022eeg] has generated results on new dataset for approaches existing on old dataset which shows that EEGNet [lawhern2018eegnet] reported about 30%, and EEG-Channel Net [palazzo2020decoding] saw a jump to approximately 50% accuracy.
All the existing approach on our task utilizes either only CNN architectures or only sequential architectures to find only local features or only sequential features from EEG signal. Our approach on contrary, uses both local as well as sequential features extracted by CNN and sequential architecture (Bi-LSTM or transformer) combinedly.
While our CNN-transformer architecture is similar to the Conformer architecture [song2022eeg], there are key differences: The Conformer architecture with its potential to first finds local temporal features and then spatial features which are then averaged before giving it to the transformer. This averaging is very well suitable for some well established EEG tasks such as motor imagery and emotion recognition paradigms, where the discrimination is relatively simpler with high accuracy in literature, whereas in our CNN-transformer which uses windows of EEG signal in its approach, averaging will result in loss of information. Therefore, the the first layer of convolution does noise removal and local temporal and spatial features are calculated by further layers for visual stimulus recognition, arguably a harder problem.
Our architecture is much simpler than complex architectures like Vision Transformers (ViT) [han2022survey]. While the concept of learning local embeddings followed by a transformer model is similar, ViT first split the image into fixed-size patches and then linearly embed these patches which are given to encoder of the Transformer. While in our CNN-Transformer approach, the 1D CNN layers learn local spatio-temporal contexts. On the other hand, ViT architecture flatten the patches to form sequence. While CNN-Transformer finds local temporal features of each channel to form a sequence. The architecture identifies sequential features, where each sequence represents local temporal features, and each local temporal feature contains spatial information across all channels.
@@ 4 Methodology
Figure 2: The architecture diagram shows the following steps: An EEG sample is divided into $11$ segments of $220$ samples each during pre-processing. This is followed by a Feature Extraction block which uses a CNN to capture spatial and temporal information. then a Sequential block, which can be a Bi-LSTM or Transformer, learns sequential relations. Finally, a Classifier block with dense layers to classify the embeddings
@@ 4.2 Feature Extraction Block
This block contains CNN to separately extract spatial and temporal features across EEG channels and time axis. It performs 1-D convolution for three layers. 1-D convolution not only contains less parameters as compared to 2-D convolution but also helps in extracting temporal features from each channel and spatial features across channels, explicitly. This provides more control over generating embeddings in temporal and channel information without intermixing them. This is well explained in [mishra2021eeg] and [mishra2020analyzing]. The first two layers perform convolution along time (activation function = Relu, kernel size = (1,35) and stride = (1,2)) while last layer do convolution over channel (activation function = sigmoid, kernel size = (128,1) and stride = (1,1)). The convolution layers over time find features across different time stamps while convolution layer over channel finds features across all channels for each extracted feature of time stamp. Each convolution layer has 25 filters. The final output of feature extractor is reshaped to (30,25) and fed to sequential block.
@@ 4.3 Bi-LSTM and Transformer Blocks
EEG signal can be fed to LSsTM in two ways: either channel[spampinato2017deep] wise or all EEG channel vector per time step. In common LSTM architecture whole EEG signal per time step is fed to LSTM. In Stacked LSTM, multiple layer of LSTM are used. Output of first layer is given to next layer. The final output of last layer at the last time step is the encoding of EEG signal[mishra2022eeg], which can further be used for classification. In the Bi-LSTM block, we have stacked Bi-LSTM to capture information in both forward and backward direction.
Figure 3: Architecture of stacked Bi-LSTM in sequential block which is fed with extracted features in both forward and reverse manner in order to learn sequence in both forward and backward direction. Present feature is represented as $x_{t}$ while previous feature is represented as $x_{t-1}$.
@@
The Transformer model, consisting of an encoder and decoder architecture for sequence to sequence mapping. The encoder processes the input data utilizing self-attention mechanisms to weigh the non-local importance among different parts of data. This processed information is then passed to the decoder, which also employs self-attention and additionally uses what is known as cross-attention to focus on relevant parts of the input sequence when generating new output sequence. This allows Transformers to efficiently handle sequences of data, capturing complex relationships within data compared to recurrent layers.
The Architecture mentioned in [vaswani2017attention] consist of encoder and decoder for sequence to sequence mapping. As our take only involves EEG sequence classification, we therefore use only the encoder part of the transformer. The encoder of the transformer for our approach finds non-local relations within the sequential feature embeddings provided by the CNN using self-attention. The proposed architecture as shown in Figure 4 has 1 encoder layer with 8 heads for multi-head attention and 2 dense layers in feed forward network.
@@ 6.1 Experimental Setup
For our experiment, the split of 80% training, 10% validation and 10% test data is used as done in [tao2021gated]. Each split contains data of each class for each subject. EEG data of all subjects for single image of a particular class is also taken into single split (either training or validation or testing) for uniformity. While splitting data it was observed that 33rd class has irregular data among different subjects. To maintain regularity among subject and class data, the 33rd class is removed, this resulting in a total of 39 classes.
@@ 6.2 Results
The result of the architectures and comparison with other existing work is shown in Table 1. These are either from their respective papers or from [mishra2022eeg], which reports a comparison of various approaches.
@@
Approach Accuracy (%)
Stacked Bi-LSTM [fares2019eeg] 22
EEGCVPR40[spampinato2017deep] 26
SyncNet [li2017targeting] 27
EEGNet [lawhern2018eegnet] 32
EEG-ChannelNet [palazzo2020decoding] 36
GRU Gated Transformer [tao2021gated] 46
SigTanh Gated Transformer [tao2021gated] 49
EfficientNet + SVM [mishra2022eeg] 64
Proposed Bi-LSTM approach 71
Proposed Transformer approach 59
We observe that our CNN-Bi-LSTM based method outperforms all existing approaches and the CNN-Transformer based method also outperform the most and better than all existing Transformer based approaches. Both our approaches have taken the leverage of both local as well as sequential features to perform the task while most existing approaches are limited to either only local features or sequential features only. The Bi-LSTM approach even performs better over approach in [zhang2020bi] which involves converting EEG signals to gray-scale image.
Table 2: Performance metrics for the proposed Bi-LSTM model
Accuracy Recall Precision F1-Score
71% 71% 72% 71%
We note that, presently the CNN-Transformer methods (ours as well as other methods) have low performance compared to the CNN-Bi-LSTM approach. We believe that a reason for this could be due to smaller window length, since Transformers usually expects long duration data to better learn non-local dependencies. Moreover, the architecture and training dynamics of Transformers might necessitate adjustments for optimal performance on EEG datasets. Factors such as the choice of hyper parameters, the design of the attention mechanism, and the amount of training data available could significantly impact the effectiveness of Transformers as this is an unconventional application domain. Hence, the unique challenges posed by EEG data require bespoke modifications to the standard Transformer model to fully leverage its capabilities.
Having said that, from Table 1, we also note that the Transformer based framework in this work outperforms the other existing Transformer based approaches which uses a gated network (using either GRU Gate or SigTanh Gate) in place of the residual network of the encoder of Transformer. Thus, this work highlights the potential of using Transformer based architectures in this domain.
Figure 5: Bar graph for accuracy of individual classes
Table 2 and Figure 5 also provide further results for our CNN-Bi-LSTM model which shows that the performance is consistently distributed for most classes, and the method achieve high scores for precision, recall and F1-score, which highlights the recognition of individual classes.
@@ 6.3 Analysing embeddings and brain mappings
Typically, in many works involving neuroimage/signal analysis, such topographical maps are used to highlight differences between conditions. In this case the conditions correspond to the classes. While the t-SNE visualizations demonstrate the difference in the class embeddings from the neural networks, the topographical maps serve a complementary purpose of suggesting, that there are also differences in the spatial distribution of brain activity across the classes. When associated with t-SNE plots, this leads to an insightful hypothesis that such original differences in the signal patterns can get encoded in the network embeddings. This additional layer of analysis would contribute to the explainability of such methods by linking brain activation patterns to the classified visual stimuli.
@@ 7.0.1 Acknowledgements
[…]